OpenStudy (anonymous):

HELP!!!!!!!!! Given the following endpoint (A) and midpoint (M), find the other endpoint (B). M(2, 3) A(0, -1)

OpenStudy (anonymous):

You know that M is the midpoint between A and the other point, so that means it is halfway there. How much does the x and y change from the endpoint to the midpoint?

OpenStudy (anonymous):

2,2?

OpenStudy (anonymous):

Look at the individual x and y parts of the ordered pairs: (2,3), x=2 y=3 (0,-1), x=0 y=-1 How much does x change by and how much does y change by between these points?

OpenStudy (anonymous):

I dont get it . do i supposed to subtract or what ?

OpenStudy (anonymous):

You find the difference between each of the x's and each of the y's. So yes you'd find the difference.

OpenStudy (anonymous):

2-0 = 2 3-(-1) = 2

OpenStudy (anonymous):

3-(-1) is actually 4, be careful with that

OpenStudy (anonymous):

so it will be (2,4)

OpenStudy (anonymous):

But yes now you found the change in x and the change in y. Since that is the change from the beginning to the midpoint, the other endpoint will be twice as far. Could you tell me what you think needs to be done now?

OpenStudy (anonymous):

multiply by 2 ?

OpenStudy (anonymous):

And the 2 and 4 are actually just vector quantities, don't worry about what that means yet but it's not used to make a point.

OpenStudy (anonymous):

And yes, you multiply by 2. What new point does that give us?

OpenStudy (anonymous):

4,8 ?

OpenStudy (anonymous):

No, multiply the 2 and 4 by 2, and then use that new change in x and y to find the new point

OpenStudy (anonymous):

4 is the change in x and 8 is the change in y from the initial starting point, so what's the point on the other end?

OpenStudy (anonymous):

4,7 ?

OpenStudy (anonymous):

Im really confused

OpenStudy (anonymous):

Yep! That point is twice as far as the midpoint correct?

OpenStudy (anonymous):

Because the x and y changes by twice as much

OpenStudy (anonymous):

Ok I get it. THANK YOU!!! :)

OpenStudy (anonymous):

Haha I'm glad you understand now ^^